OpenStudy (anonymous):

Find the GCF. 12 and 45 A. 45 B. 3 C. 12 D. 180

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

I will do it if you help me with my question.

OpenStudy (misty1212):

lol not not C it is almost always c but not this time

OpenStudy (anonymous):

B

OpenStudy (anonymous):

yes it is your thinking of LCD its GCF and the GCF of 12 and 45 is 3

OpenStudy (anonymous):

Can someone help me with my question? Please? I really need help!

OpenStudy (anonymous):

one sec @Aelia6th

OpenStudy (anonymous):

So is it C or B guys because now im super confused.

OpenStudy (misty1212):

?? is 12 a factor of 45? no of course not so 12 cannot be the greatest common factor of 12 and 45

OpenStudy (misty1212):

\[12=2\times 2\times 3\] \[45=5\times 3\times 3\] the only factor they have in common is \(3\)

OpenStudy (misty1212):

it is the only common factor, so it is also the "greatest common factor"

OpenStudy (misty1212):

i have no idea where the 12 comes from in the answer they just put that answer choice in there to confuse you because 12 is one of the two numbers

OpenStudy (mathstudent55):

Prime factorization of 12: \(2^2 \times 3\) Prime factorization of 45: \(3^3 \times 5\) GCF: use only common factors with smaller exponent. 2 is not a common factor (only in 12). 5 is not a common factor (only in 45). 3 is common. Use 3 (lower exponent), not 3^2 (higher exponent). GCF of 12 and 45: 3

OpenStudy (anonymous):

c

OpenStudy (mathstudent55):

Answer is B

OpenStudy (anonymous):

The answer is B becuase the equal multiplication is 3 between 12 and 45

OpenStudy (anonymous):

It's B. The GCF, is the largest number that divided both. It can't be A, since 45 doesn't divide 12. It can't be C, since 12 doesn't divide 45. (well it does, just not into a whole number) And it can't D, since they asking for the common factor, not the number they both can nultiply into...

OpenStudy (anonymous):

12/3 =4 45/3 = 15